Question 1
Suppose the representative consumer’s utility depends on goods, c and leisure , 24−l, where l is the quantity of hours worked: u = ln c+(24−l).
And suppose the consumer’s production function is c = 6 l. (a) find the equilibrium c, l and u. [24 Marks]
(b) What changes in the production over time would be required to gen- erate economic growth? [4 Marks]
1
Expert's answer
2019-08-15T09:22:25-0400

a) u = ln c+(24−l), I=24−l, c = 6 l

equilibrium I=24, c=144, u=24n(144)

b) increase in leisure and decrease quantity of hours work